# Exploring Variance Reduction in Importance Sampling for Efficient DNN Training

### v1

@@ 7.3 DNN model, training, and evaluation settings
The DNN models and training configurations for each dataset are summarized in Table 1. For each dataset, multiple experiments were conducted by varying the total number of training iterations (“Total-iters” in the table). For CINIC-10, two different weight-decay settings were evaluated. The minibatch size was fixed at 128 as a common setting across all datasets. The learning rate was initialized with the values specified in Table 1 and decayed using cosine scheduling, ensuring the learning rate reaches zero at the specified total number of training iterations.
@@
LR: learning rate, WD: weight decay, DA: data augmentation, SD: stochastic depth [9],
CE: cross entropy, BCE: binary cross entropy
@@ 7.4.1 Comparison of EMAIS and SGD-Uni with dynamic minibatch size
Figure 2 presents the results, including the training loss and test error for each model, as well as the transitions of $N_{\mathrm{ems}}$ recorded during EMAIS training. For comparison, we also show the training loss and test error of SGD-Uni with a fixed minibatch size of $N=128$. The figure shows that both the training loss and test error follow nearly the same trend for EMAIS with a fixed minibatch size of $N=128$ and SGD-Uni with a dynamic minibatch size of $N=N_{\mathrm{ems}}$. This suggests that training with importance sampling, given an appropriate $W$, effectively mimics the effect of increasing the minibatch size. Moreover, $N_{\mathrm{ems}}$ serves as an effective estimator of the extent to which the minibatch size is adjusted by importance sampling. The figure also illustrates the superior performance of EMAIS compared with SGD-Uni under the condition that the minibatch size is fixed at $N=128$. As described in subsection 6.2, the reason $N_{\mathrm{ems}}$ remains flat in the early stages of training is that EMAIS estimates the initial values of the moving statistics during this period, without applying importance sampling.
@@ 7.4.2 Comparison of prediction accuracy
140000 $13.49(\pm 0.08)$ $13.55(\pm 0.11)$ $13.31(\pm 0.09)$ $13.80(\pm 0.10)$ $13.79(\pm 0.13)$ $13.52(\pm 0.11)$ $13.66(\pm 0.09)$ $\textbf{13.13}(\pm 0.03)$
1e-5 17500 $\textbf{15.27}(\pm 0.06)$ $15.52(\pm 0.13)$ $\textbf{15.23}(\pm 0.03)$ $\textbf{15.26}(\pm 0.11)$ $15.61(\pm 0.11)$ $15.51(\pm 0.07)$ $16.18(\pm 0.25)$ $\textbf{15.19}(\pm 0.15)$
@@
Table 4: mAP on ChestX-ray14 test set ($\uparrow$)
@@ 7.4.3 Comparison of training time
The training times for DNN models on FMNIST, CINIC-10, and ChestX-ray14 are summarized in Table 5, Table 6, and Table 7, respectively. [^3] The experiments were conducted on a machine equipped with an $\text{Intel}^{\text{\textregistered}}$ $\text{Xeon}^{\text{\textregistered}}$ Gold 6134 CPU and $\text{NVIDIA}^{\text{\textregistered}}$ $\text{Tesla}^{\text{\textregistered}}$ V100 GPU. Each table reports the mean training time in seconds, with the numbers in parentheses representing the standard deviations. The time required for DNN training can vary due to factors beyond the computational complexity of the learning algorithm, such as implementation details or the impact of other computational tasks when using shared computing resources. Therefore, while the current results are not suitable for a detailed comparison of computational complexity, they do provide an approximate understanding of the overall trends in training time.
Table 5: Training time (s) for FMNIST
Table 6: Training time (s) for CINIC-10
1e-4 17500 $1169.3(\pm 76.4)$ $1111.3(\pm 13.7)$ $1188.2(\pm 65.4)$ $3496.6(\pm 139.2)$ $1116.2(\pm 30.7)$ $1167.5(\pm 29.6)$ $1137.5(\pm 33.2)$ $1153.6(\pm 47.7)$
35000 $2329.3(\pm 167.3)$ $2322.7(\pm 151.5)$ $2327.2(\pm 50.1)$ $7140.6(\pm 397.3)$ $2250.7(\pm 68.3)$ $2342.5(\pm 66.6)$ $2272.0(\pm 39.9)$ $2280.6(\pm 87.5)$
70000 $4693.2(\pm 168.4)$ $4498.9(\pm 27.3)$ $4552.2(\pm 38.5)$ $13596.8(\pm 287.9)$ $4416.5(\pm 115.1)$ $4591.8(\pm 49.6)$ $4601.4(\pm 91.5)$ $4546.3(\pm 88.0)$
1e-5 17500 $1164.0(\pm 54.2)$ $1174.8(\pm 63.4)$ $1183.8(\pm 61.2)$ $3443.2(\pm 120.3)$ $1114.6(\pm 39.3)$ $1185.3(\pm 27.3)$ $1112.9(\pm 21.3)$ $1237.2(\pm 29.2)$
35000 $2250.6(\pm 50.7)$ $2415.6(\pm 163.2)$ $2307.8(\pm 43.3)$ $6885.3(\pm 211.3)$ $2219.1(\pm 63.6)$ $2364.7(\pm 58.1)$ $2254.2(\pm 36.4)$ $2473.4(\pm 37.8)$
Table 7: Training time (s) for ChestX-ray14
17500 $7768.5(\pm 14.2)$ $7758.2(\pm 2.0)$ $7866.1(\pm 43.1)$ $23824.4(\pm 25.7)$ $7759.7(\pm 3.1)$ $7773.8(\pm 3.1)$
@@
As shown in Tables 5, 6 and 7, there were no significant differences in training time among the methods, except for Presampling-IS. For Presampling-IS, the computational overhead was notably large because it requires performing calculations on large minibatches beforehand, which resulted in the observed differences in training time. Taking into account the results from the previous section and these findings, EMAIS can train more accurate models with approximately the same computational time as the widely used SGD-Scan.
@@ 8 Conclusion
We proposed a method for estimating the variance reduction of loss gradient estimators in DNN training using importance sampling, compared to uniform sampling. Since the proposed method requires only minibatches generated through importance sampling, it enables the online estimation of the variance reduction rate during DNN training. By leveraging the proposed method, we also proposed an EMS to enable automatic learning rate adjustment. We developed an absolute metric to evaluate the efficiency of importance sampling and designed an algorithm to estimate importance score using moving statistics. Through numerical experiments, we demonstrated that the proposed method consistently achieved higher accuracy than uniform sampling, while maintaining comparable computational time. We also showed that the proposed method outperformed related importance-sampling methods, highlighting its effectiveness and efficiency.
@@ Appendix A Proofs and formula derivations
A.1 Proof for expectation equivalence
$\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$ $=\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\nabla_{\theta}\mathcal{L}(\theta;i)\right]$
$=\sum_{i=1}^{M}p_{\mathrm{unif}}(i)\nabla_{\theta}\mathcal{L}\left(\theta;i\right)$
$=\sum_{i=1}^{M}p_{\mathrm{is}}(i;W)\frac{p_{\mathrm{unif}}(i)}{p_{\mathrm{is}}(i;W)}\nabla_{\theta}\mathcal{L}\left(\theta;i\right)$
$=\sum_{i=1}^{M}p_{\mathrm{is}}(i;W)r\left(i;W\right)\nabla_{\theta}\mathcal{L}\left(\theta;i\right)$
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\nabla_{\theta}\mathcal{L}(\theta;i)\right]$
$=\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right].$
@@ A.1 Proof for expectation equivalence
$=\sum_{\theta_{k}\in\theta}\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right]$
$=\sum_{\theta_{k}\in\theta}\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[\left(r(i;W)\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[\sum_{\theta_{k}\in\theta}\left(r(i;W)\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
@@
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{\theta_{k}\in\theta}\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right]$
$=\sum_{\theta_{k}\in\theta}\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\left(\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
@@
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|^{2}\right]-\left\|\mu\right\|^{2}.$
@@
$=\left(\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|\right]\right)^{2}-\left\|\mu\right\|^{2}.$

### v2

@@ 9.3 DNN model, training, and evaluation settings
The DNN models and training configurations for each dataset are summarized in Table 1. Specifically, LeNet5 [20], ResNet18 [9], and ResNet50 [9] are employed for the FMNIST, CINIC-10, and ChestX-ray14 datasets, respectively. For each dataset, multiple experiments were conducted by varying the total number of training iterations (“Total-iters” in Table 1). For CINIC-10, two different weight-decay settings were evaluated. The minibatch size was fixed at 128 as a common setting across all datasets. For all methods, the learning rate was initialized with the values specified in Table 1 and decayed using cosine scheduling, ensuring that it reaches zero at the specified total number of training iterations.
@@
LR: initial learning rate (decayed using cosine scheduling), WD: weight decay, DA: data augmentation,
SD: stochastic depth [10], CE: cross entropy, BCE: binary cross entropy
@@ 9.4.1 Comparison of EMAIS and SGD-Uni with dynamic minibatch size
Figure 2 presents the results, including the training loss and test error for each model, as well as the transitions of $N_{\mathrm{ems}}$ recorded during EMAIS training. For comparison, the training loss and test error of SGD-Uni and SGD-Scan with a fixed mini-batch size of $N=128$ are also shown. The figure shows that both the training loss and test error follow nearly the same trend for EMAIS with a fixed minibatch size of $N=128$ and SGD-Uni with a dynamic minibatch size of $N=N_{\mathrm{ems}}$. This suggests that training with importance sampling, given an appropriate $W$, effectively mimics the effect of increasing the minibatch size. Moreover, $N_{\mathrm{ems}}$ serves as an effective estimator of the extent to which the minibatch size is adjusted by importance sampling. The figure also illustrates the superior performance of EMAIS compared with SGD-Uni and SGD-Scan under the condition that the minibatch size is fixed at $N=128$. As described in subsection 6.2, the reason $N_{\mathrm{ems}}$ remains flat in the early stages of training is that EMAIS estimates the initial values of the moving statistics during this period, without applying importance sampling.
@@ 9.4.2 Comparison of prediction accuracy
[…]
1e-5 17500 $\textbf{15.27}(\pm 0.06)$ $15.52(\pm 0.13)$ $\textbf{15.23}(\pm 0.03)$ $\textbf{15.26}(\pm 0.11)$ $61(\pm 0.11) 15.51$(±0.07)15.51($\pm$ 0.07) $16.18(\pm 0.25)$ $\textbf{15.19}(\pm 0.15)$
@@
Table 4: Mean average precision (mAP) on ChestX-ray14 test set ($\uparrow$)
@@ 9.4.3 Comparison of training time
The training times for DNN models on FMNIST, CINIC-10, and ChestX-ray14 are summarized in Table 5.333The experiments were conducted on a machine equipped with an $\text{Intel}^{\text{\textregistered}}$ $\text{Xeon}^{\text{\textregistered}}$ Gold 6134 CPU and an $\text{NVIDIA}^{\text{\textregistered}}$ $\text{Tesla}^{\text{\textregistered}}$ V100 GPU. The table reports the average training time per iteration for each dataset. [^4] For detailed training times under each training configuration, please refer to subsection C.3. The time required for DNN training can vary due to factors beyond the computational complexity of the learning algorithm, such as implementation details or the impact of other computational tasks when using shared computing resources. Therefore, while the current results are not suitable for a detailed comparison of computational complexity, they do provide an approximate understanding of the overall trends in training time.
Table 5: Per-iteration average training time (in seconds)
Dataset SGD-scan SGD-unif RAIS Presampling-IS Confidence ConfVar Self-paced EMAIS
FMNIST $8.60\times 10^{-3}$ $8.62\times 10^{-3}$ $1.08\times 10^{-2}$ $5.22\times 10^{-2}$ $6.21\times 10^{-3}$ $9.13\times 10^{-3}$ $8.66\times 10^{-3}$ $7.32\times 10^{-3}$
CINIC-10 $62\times 10^{-2} 6.6$$3\times 10^{-2} 6.6$$63\times 10^{-2}$ $1.98\times 10^{-1}$ $6.35\times 10^{-2}$ $6.70\times 10^{-2}$ $50\times 10^{-2} 6.6$$67\times 10^{-2}$
ChestX-ray14 $45\times 10^{-1} 4.4$$45\times 10^{-1}$ $4.50\times 10^{-1}$ $1.36\times 10^{0}$ $45\times 10^{-1} 4.4$$46\times 10^{-1}$
@@
As shown in Table 5, the per-iteration training time differs across datasets, which is partly due to the size of the DNN models employed for each dataset. When focusing on individual datasets, we observe no significant differences in training time among the methods, except for Presampling-IS. For Presampling-IS, the computational overhead was notably large because it requires performing calculations on large minibatches beforehand, which resulted in the observed differences in training time. Taking into account the results from the previous section and these findings, EMAIS can train more accurate models with approximately the same computational time as the widely used SGD-Scan.
9.5 Evaluation with the Adam optimizer
The above evaluation was conducted using SGD as the optimizer. In this section, we present the evaluation results obtained using the Adam optimizer. The experimental settings largely follow those in Table 1, except for the learning rate configuration, which was modified as follows (consistently applied across all datasets): the initial learning rate was set to 0.001 and decayed according to cosine scheduling. We denote the method that samples minibatches using uniform sampling as Adam-Uni, and the one that uses uniform sampling without resampling as Adam-Scan. In addition to these, we evaluate RAIS and EMAIS using the Adam optimizer, referred to as RAIS-Adam and EMAIS-Adam, respectively. For EMAIS-Adam, we applied the learning rate adjustment based on $N_{\mathrm{ems}}$ for Adam, which is denoted as $\epsilon_{\mathrm{ems}}^{\mathrm{adam}}$ in subsection 5.2.
The results are presented in Table 6, showing the mean values of the evaluation metrics, with the numbers in parentheses indicating the standard deviations over five trials. Bold values indicate either the best result or those whose mean falls within the standard deviation of the best result. As shown in Table 6, EMAIS-Adam consistently outperforms the uniform sampling-based methods, Adam-Scan and Adam-Uni, across all datasets and training configurations. These results demonstrate the effectiveness of importance sampling over uniform sampling, even when combined with the Adam optimizer. When compared to RAIS-Adam, EMAIS-Adam shows comparable performance overall. In several configurations, such as those for CINIC-10, the two methods yield nearly identical results. However, for ChestX-ray14, EMAIS consistently outperforms RAIS across all training durations. This underscores the effectiveness of our method not only on standard classification tasks, but also on more complex problems such as medical image analysis with imbalanced multi-label data.
Table 6: Evaluation results on FMNIST, CINIC-10, and ChestX-ray14 using the Adam optimizer. For FMNIST and CINIC-10, the values represent test error (%, lower is better), while for ChestX-ray14, the values indicate mean average precision (mAP; higher is better).
Dataset WD Total-iters Adam-Scan Adam-Uni RAIS-Adam EMAIS-Adam
FMNIST ($\downarrow$) 1e-3 6250 $10.99(\pm 0.31)$ $10.96(\pm 0.14)$ $10.20(\pm 0.09)$ $\textbf{10.00}(\pm 0.13)$
12500 $9.75(\pm 0.10)$ $9.81(\pm 0.17)$ $9.18(\pm 0.22)$ $\textbf{8.92}(\pm 0.07)$
25000 $9.09(\pm 0.16)$ $9.00(\pm 0.25)$ $\textbf{8.94}(\pm 0.21)$ $\textbf{8.82}(\pm 0.17)$
CINIC-10 ($\downarrow$) 1e-4 17500 $17.46(\pm 0.08)$ $17.90(\pm 0.11)$ $\textbf{17.32}(\pm 0.11)$ $\textbf{17.43}(\pm 0.07)$
35000 $17.02(\pm 0.15)$ $17.01(\pm 0.11)$ $\textbf{16.78}(\pm 0.13)$ $\textbf{16.74}(\pm 0.09)$
70000 $16.42(\pm 0.14)$ $16.56(\pm 0.09)$ $\textbf{16.12}(\pm 0.06)$ $\textbf{16.13}(\pm 0.10)$
5e-5 17500 $18.86(\pm 0.23)$ $18.87(\pm 0.15)$ $\textbf{18.36}(\pm 0.07)$ $18.45(\pm 0.16)$
35000 $17.73(\pm 0.10)$ $17.72(\pm 0.16)$ $\textbf{17.55}(\pm 0.13)$ $\textbf{17.49}(\pm 0.15)$
70000 $17.28(\pm 0.22)$ $17.40(\pm 0.13)$ $\textbf{17.18}(\pm 0.09)$ $\textbf{17.23}(\pm 0.11)$
ChestX-ray14 ($\uparrow$) 1e-4 17500 $15.14(\pm 0.21)$ $15.19(\pm 0.15)$ $15.47(\pm 0.24)$ $\textbf{15.98}(\pm 0.10)$
35000 $16.90(\pm 0.13)$ $17.03(\pm 0.08)$ $17.10(\pm 0.24)$ $\textbf{17.59}(\pm 0.14)$
70000 $18.59(\pm 0.19)$ $18.46(\pm 0.14)$ $18.38(\pm 0.23)$ $\textbf{19.08}(\pm 0.14)$
9.6 Assessment of estimation with logit gradients
This section summarizes the results of evaluating the estimation accuracy of each metric when using gradients with respect to the logits, instead of the full parameter gradients discussed in section 7. In this analysis, we used LeNet5 trained on FMNIST as a representative example, following the training configuration with 25000 total iterations as shown in Table 1. Although LeNet5 is a relatively small convolutional neural network, it contains over 60000 trainable parameters, making its parameter space substantially higher-dimensional than that of the logits, which have a dimensionality of 10 in this case.
We first present the results of comparing the per-sample gradient norms computed using full parameter gradients with those computed using logit gradients. For models at training iterations 5000 and 25000, we computed both $\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|$ and $\left\|\nabla_{z}\tilde{\mathcal{L}}(z;i)\right\|$ for each of 500 data points $i$ randomly sampled from the training dataset. The results are presented in Figure 3. Each scatter plot compares the logit-based gradient norm (horizontal axis) with the corresponding full parameter gradient norm (vertical axis) across data points. As shown in the figure, although $\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|$ and $\left\|\nabla_{z}\tilde{\mathcal{L}}(z;i)\right\|$ differ in scale, they exhibit a clear proportional relationship. This observation is consistent with previous findings reported in [13].
Figure 3: Relationship between the full gradient and logit gradient norms for models at different training iterations.
Next, we evaluated the accuracy of trace estimates of the gradient variance based on the logit-gradient approximation, as discussed in section 7, using models at training iterations 5000 and 25000. For each model, we first computed the optimal importance weights $W^{\ast}$, and then generated multiple variants of importance weight vectors $W$ by adding controlled noise to $W^{\ast}$. Using each $W$, we constructed a single minibatch of size 128 via importance sampling. For every data point $i$ in the sampled minibatch, we computed both $\nabla_{\theta}\mathcal{L}(\theta;i)$ and $\nabla_{z}\tilde{\mathcal{L}}(z;i)$. These values were then used to estimate the trace of the gradient variance in accordance with Proposition 4.1 and Eqs. 10, 11 and 12. The results are presented in Figure 4. Each scatter plot compares the trace estimates of the gradient variance computed using logit gradients (horizontal axis) with those computed using full parameter gradients (vertical axis), across minibatches generated with different $W$. As shown in the figure, the trace estimates of the gradient variance computed using logit gradients exhibit a clear linear relationship with those computed using full parameter gradients, although the two quantities differ substantially in scale. For reference, the fitted regression lines are shown in red in each plot. While some variation is observed across models, both the estimated regression coefficients and intercepts remain within a similar order of magnitude.
Figure 4: Comparison of trace variance estimates in Proposition 4.1, computed using either full parameter gradients (plotted on the $y$-axis) or logit gradients (on the $x$-axis), as discussed in section 7, for models at training iterations 5000 and 25000. The red lines represent fitted regression lines.
Finally, we compared the estimates of $N_{\mathrm{ems}}$ and $\mathcal{S}(W)$ obtained using logit gradients and full parameter gradients, respectively. For each minibatch generated by the procedure described above, we estimated the trace of the gradient variance and subsequently computed $N_{\mathrm{ems}}$ and $\mathcal{S}(W)$ based on these trace estimates. The results are shown in Figure 5. Each scatter plot compares the estimated scores, either $N_{\mathrm{ems}}$ or $\mathcal{S}(W)$, computed using logit gradients (horizontal axis) with those computed using full parameter gradients (vertical axis), across the generated minibatches. As shown in the figure, compared to estimates based on full parameter gradients, those obtained using logit gradients tend to slightly overestimate $N_{\mathrm{ems}}$ and slightly underestimate $\mathcal{S}(W)$. Nevertheless, the two sets of estimates exhibit a strong linear relationship across all settings, indicating that the logit-based approximation can serve as a reliable substitute. Notably, in contrast to gradient norms or trace variance estimates, the values of $N_{\mathrm{ems}}$ and $\mathcal{S}(W)$ computed using logit gradients remain within the same order of magnitude as those obtained from full gradients. This property is especially beneficial when absolute magnitudes influence downstream decisions, such as learning rate adjustment. As discussed in section 7, this consistency can be attributed to the fact that both $N_{\mathrm{ems}}$ and $\mathcal{S}(W)$ are computed based on the ratio (or the ratio of differences) of trace variance estimates, which inherently normalizes the scale differences between full and logit gradients.
Figure 5: Scatter plots comparing estimates based on full parameter gradients versus logit-based approximations for (a, b) the effective minibatch size $N_{\mathrm{ems}}$ and (c, d) the efficiency metric $\mathcal{S}(W)$, for models at training iterations 5000 and 25000.
9.7 Ablation study of learning rate adjustment
We report the results of an ablation study aimed at demonstrating the effectiveness of the learning rate adjustment proposed in subsection 5.2. Using the FMNIST configuration with a total of 25000 iterations from Table 1, we trained the model with EMAIS, both with and without the learning rate adjustment. The results are shown in Figure 6, including the training loss, test error, effective minibatch size $N_{\mathrm{ems}}$, and the learning rate $\tilde{\epsilon}$ actually applied during training. For EMAIS with learning rate adjustment, $\tilde{\epsilon}=\epsilon_{\mathrm{ems}}$ is computed as in Eq. 8. For EMAIS without adjustment, $\tilde{\epsilon}=\epsilon$, which corresponds to the original learning rate with cosine decay. The figure also includes the scaled learning rate, defined as $\frac{N}{N_{\mathrm{ems}}}\tilde{\epsilon}$, which accounts for the virtual increase in minibatch size from $N$ to $N_{\mathrm{ems}}$. From the figure, we observe that, under the learning-rate-adjustment setting, the learning rate is scaled up with increasing $N_{ems}$, such that the scaled learning rate follows the cosine decay schedule. By contrast, in the absence of learning-rate adjustment, although the actual learning rate used during training follows cosine decay, the scaled learning rate effectively applies a smaller rate than the decay schedule. Furthermore, this reduced scaled learning rate slows the decrease of the training loss and thereby degrades the test error reduction rate.
Figure 6: Ablation study of learning rate adjustment on FMNIST. From top to bottom, the panels show: (1) training loss (log scale); (2) test error (log scale); (3) effective minibatch size $N_{\mathrm{ems}}$; (4) learning rate $\tilde{\epsilon}$ actually applied during training, which is given by $\tilde{\epsilon}=\epsilon_{\mathrm{ems}}$ for EMAIS with learning rate adjustment, and $\tilde{\epsilon}=\epsilon$ for EMAIS without learning rate adjustment; and (5) scaled learning rate that accounts for the virtual increase in minibatch size from $N$ to $N_{\mathrm{ems}}$, given by $\frac{N}{N_{\mathrm{ems}}}\tilde{\epsilon}$.
9.8 Effect of additional updates to internal states in EMAIS
In the proposed method, the importance sampling weights are computed based on a moving average of per-sample gradient norms, which are maintained as internal states. However, these internal states are updated only for the samples included in the mini-batches during training. In this study, we experimentally investigate whether training can be accelerated by periodically computing the gradient norms of samples that are not included in the current mini-batch and updating their corresponding importance weights. As representative examples, we used the FMNIST and CINIC-10 configurations with total-iters $25000$ and $70000$, respectively, as listed in Table 1.
In addition to the internal state updates based on mini-batches, we performed supplementary updates to the internal states using the following procedure. For FMNIST, every $500$ training iterations, and for CINIC-10, every $1000$ iterations, we froze the model parameters and computed the per-sample gradient norms for the entire training set. These values were then used to invoke the UpdateStats procedure from Algorithm 1 to update the internal states. Note that these gradient computations were not used to update the model parameters. A comparison of the training loss, test error, and the efficiency score $\mathcal{S}(W)$ is presented in Figure 7. As shown in the figure, the efficiency score $\mathcal{S}(W)$ tends to be slightly lower when additional state updates are applied, particularly during the early stages of training, compared to the original EMAIS. However, the training loss and test error remain largely unchanged. These findings suggest that, in the proposed algorithm, estimating importance sampling weights solely based on mini-batches during training—which entails lower computational overhead—has only a limited effect on overall training performance.
Figure 7: Performance of EMAIS with periodic additional updates of internal state variables on FMNIST and CINIC-10.
@@ 10 Conclusion
We proposed a method for estimating the variance reduction of loss gradient estimators in DNN training using importance sampling, compared to uniform sampling. Since the proposed method requires only minibatches generated through importance sampling, it enables the online estimation of the variance reduction rate during DNN training. By leveraging the proposed method, we also proposed an EMS to enable automatic learning rate adjustment. We developed an absolute metric to evaluate the efficiency of importance sampling and designed an algorithm to estimate importance weight using moving statistics. Through numerical experiments, we demonstrated that the proposed method consistently achieved higher accuracy than uniform sampling, while maintaining comparable computational time. We also showed that the proposed method outperformed related importance-sampling methods, highlighting its effectiveness and efficiency.
As a potential extension, it would be worth exploring the application of our method to distributed learning. Alain et al. [1] proposed a distributed importance sampling approach in which, at every training iteration, the gradient norm for each training sample is computed across workers, aggregated on a master node, and then used for importance sampling and model updates. In contrast, our method estimates importance weights based on moving statistics of gradient norms, which may reduce the need for communication at every iteration. Furthermore, under a fixed data-sharding regime, each node may only need to communicate the scalar sum of its local weight vector $W$ to reconstruct the global sampling distribution and perform importance sampling entirely on-node, potentially reducing communication costs even further.
@@ Appendix A Proofs and formula derivations
A.1 Proof of Proposition 3.1
For any function $g(i)$ defined over $i\in\llbracket M\rrbracket$, the following holds:
$\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[g(i)\right]$ $=\sum_{i=1}^{M}p_{\mathrm{unif}}(i)g(i)$ (13)
$=\sum_{i=1}^{M}p_{\mathrm{is}}(i;W)\frac{p_{\mathrm{unif}}(i)}{p_{\mathrm{is}}(i;W)}g(i)$
$=\sum_{i=1}^{M}p_{\mathrm{is}}(i;W)r\left(i;W\right)g(i)$
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)g(i)\right].$
Therefore, by setting $g(i)=\nabla_{\theta}\mathcal{L}(\theta;i)$, we obtain
$\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\nabla_{\theta}\mathcal{L}(\theta;i)\right]=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\nabla_{\theta}\mathcal{L}(\theta;i)\right],$
which indicates that $\mathbb{E}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]=\mathbb{E}_{\mathrm{is}(W)}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]$.
@@
$=\sum_{k=1}^{K}\mathbb{V}_{\mathrm{is}(W)}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right]$
$=\sum_{k=1}^{K}\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[\left(r(i;W)\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[\sum_{k=1}^{K}\left(r(i;W)\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
@@
$\operatorname{tr}\left(\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta}\mathcal{L}(\theta)\right]\right)$ $=\sum_{k=1}^{K}\mathbb{V}_{\mathrm{unif}}\left[\nabla_{\theta_{k}}\mathcal{L}(\theta)\right]$
$=\sum_{k=1}^{K}\mathbb{E}_{i\sim p_{\mathrm{unif}}(i)}\left[\left(\nabla_{\theta_{k}}\mathcal{L}(\theta;i)-\mu_{k}\right)^{2}\right]$
@@
$=\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|^{2}\right]-\left\|\mu\right\|^{2},$
where the final transformation follows from Eq. 13 with $g(i)=\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|^{2}$.
@@
$=\left(\mathbb{E}_{i\sim p_{\mathrm{is}}(i;W)}\left[r(i;W)\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|\right]\right)^{2}-\left\|\mu\right\|^{2},$
where the transformation follows from Eq. 13 with $g(i)=\left\|\nabla_{\theta}\mathcal{L}(\theta;i)\right\|$.